Given a total cost function, 

TC(Q) = Q

2 + 10Q + 100

where Q represents quantity of output produced. 

(i) Find the expressions for the variable cost, fixed cost, average cost, average variable cost, 

and average fixed cost. 5

(ii) At what output level (Q) is the average cost lowest? Also find the minimum average cost.


1
Expert's answer
2021-04-26T19:21:30-0400

"TC = Q^2 + 10Q +100"

1) "TC = TVC + TFC;" "TVC = Q^2 +10Q;" "TFC = 100";

"AC = TC/Q = Q+10 +\frac{100}{Q}"

"AVC= TVC\/Q = Q+10"

2)"AC' = 1 - 100\times \frac{1}{Q^2} =0" => "\frac{1}{Q^2} = \frac{1}{100} => Q = 10"

"AC_{min} = 10 + 10 + \frac{100}{10}=30"